# It’s Not the Target, It’s the Background: Rethinking Infrared Small Target Detection via Deep Patch-Free Low-Rank Representations

### v2

@@ Abstract
Infrared small target detection (IRSTD) remains a long-standing challenge in complex backgrounds due to low signal-to-clutter ratios (SCR), diverse target morphologies, and the absence of distinctive visual cues. While recent deep learning approaches aim to learn discriminative representations, the intrinsic variability and weak priors of small targets often lead to unstable performance. In this paper, we propose a novel end-to-end IRSTD framework, termed LRRNet, which leverages the low-rank property of infrared image backgrounds. Inspired by the physical compressibility of cluttered scenes, our approach adopts a compression–reconstruction–subtraction (CRS) paradigm to directly model structure-aware low-rank background representations in the image domain, without relying on patch-based processing or explicit matrix decomposition. To the best of our knowledge, this is the first work to directly learn low-rank background structures using deep neural networks in an end-to-end manner. Extensive experiments on multiple public datasets demonstrate that LRRNet outperforms 38 state-of-the-art methods in terms of detection accuracy, robustness, and computational efficiency. Remarkably, it achieves real-time performance with an average speed of 82.34 FPS. Evaluations on the challenging NoisySIRST dataset further confirm the model’s resilience to sensor noise. The source code will be made publicly available upon acceptance.
Index Terms:
Infrared small target, low-rank matrix recovery, sparse representation, image segmentation, deep neural network.
@@ I Introduction
The remainder of this paper is organized as follows: In Section II, related works are briefly reviewed. In Section III, we present the proposed method in detail. In Section IV, the experimental results are given and discussed. Conclusions are drawn in Section V.
@@ II-C Learned Image Compression Model
III Proposed Method
III-A Overview
@@
III-B Compression and Reconstruction Module
@@
Overview of the Proposed Compression and Reconstruction Module. The proposed compression-reconstruction module can be viewed as a specialized mapping $\mathcal{F}:\mathbf{F}_{I}\rightarrow\mathbf{F}_{B}$, where $\mathbf{F}_{I}\in\mathbb{R}^{W\times H\times C}$ denotes the feature map extracted from the original image, and $\mathbf{F}_{B}$ represents the corresponding reconstructed background feature map. Notably, $\mathbf{F}_{B}$ shares the same spatial and channel dimensions as $\mathbf{F}_{I}$. Given an input feature map, we first apply a $3\times 3$ convolution to suppress local noise. The processed features are then passed through successive stages of the densely connected U-Net. At the lowest resolution level, a self-attention mechanism [4] is employed to capture long-range dependencies between features.
@@ III-D Loss Function
IV Experiment
IV-A Experimental Setup
IV-A1 Datasets
@@ IV-A2 Evaluation Metrics
where $N$ is the total number of samples, $T$, $P$, $FP$ and $TP$ denote the number of ground truth, predicted positive pixels, the number of false positive, and the number of true positive respectively.
IV-A3 Implementation Details
@@
IV-B Comparison with State-of-the-Arts
@@
IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow
@@
[…]
@@
[…]
@@ IV-B1 Quantitative Evaluation
It is worth noting that our LRRNet achieves a throughput of 82.34 FPS on an NVIDIA 3080Ti GPU with an input resolution of $512\times 512$ and a batch size of 1. In contrast, methods that rely on learning object-specific features often suffer from significantly slower inference speeds [19]. For example, under the same conditions, DNANet [11] achieves only 4.56 FPS. This performance gap is primarily due to the use of complex module designs aimed at incorporating small object information into deeper layers, which substantially increases computational overhead and slows down the model.
@@ IV-B2 Qualitative Evaluation
IV-C Ablation Study
IV-C1 The Effect of Densely Connection
@@
As shown in Table III, the densely connected U-Net significantly outperforms the baseline: IoU increases by 1.90, nIoU by 1.8, $P_{d}$ by 1.68, and $F_{a}$ decreases by 11.29. This improvement stems from the limitation of standard skip connections, which, despite preserving information, suffer from feature confusion in deeper layers [43], hindering accurate background reconstruction. Moreover, deep equilibrium models [79] suggest that even identical modules can respond differently to various features. Dense long-range connections facilitate finer-grained feature propagation, enabling better preservation of subtle structures in cluttered scenes.
TABLE III: Ablation Study on Compression–Reconstruction Module Using the IRSTD-1K Dataset.
@@
IV-C2 The Effect of Long Skip Connections Without Attention
@@
In contrast, our design intentionally omits attention mechanisms along the skip paths. As shown in Table III, this simple yet effective choice yields superior performance with statistically significant gains. We attribute this to the tendency of attention modules to overemphasize dominant structures, which, while beneficial for target enhancement, can hinder accurate low-rank background reconstruction. By avoiding such interference, our approach enables more precise modeling of the background and clearer separation between targets and clutter.
IV-C3 The Effect of Self-Attention
@@
As shown in Table III, introducing a single self-attention block significantly improves performance. However, adding multiple blocks leads to degradation. This can be attributed to the lack of inductive bias in self-attention [4]; when overused, it behaves like deep clustering—useful for compression but detrimental to reconstruction fidelity. Thus, a restrained application of self-attention is key to balancing compression efficiency with accurate reconstruction.
TABLE IV: Ablation Study on Subtraction Module Using the IRSTD-1K Dataset.
@@
IV-C4 The Effect of Subtraction Strategy
The subtraction strategy separates target-related information by subtracting the reconstructed background feature map from the corresponding feature map of the original image. If replaced with addition, the model effectively reverts to a conventional target feature learning approach. As shown in Tab. IV, the subtraction strategy significantly enhances model performance.
IV-C5 The Effect of Reconstruction Strategy
The reconstruction strategy is also a key component of LRRNet. On one hand, it guides the model to reconstruct the background; on the other, the use of MSE loss enhances the model’s robustness to Gaussian noise. As shown in Tab. IV, the reconstruction strategy has a significant impact on overall performance.
IV-C6 The Effect of ResBlock
The ResBlock plays an important role in effectively suppressing noise in the input image [55]. As shown in Tab. IV, simply replacing it with a $3\times 3$ convolution results in a significant performance drop. However, increasing the number of ResBlocks excessively can also be detrimental. Although short skip connections help preserve features, combining them through nonlinear operations may introduce feature confusion. Given the small size of the targets, even minor disturbances can lead to noticeable performance degradation.
IV-C7 The Effect of Pre-act ResBlock
The structure of the ResBlock also has a considerable impact on performance. In particular, the Pre-activation ResBlock demonstrates stronger robustness to noise compared to other variants [44], which is critical for both the compression–reconstruction and subtraction processes. As shown in Tab. III and IV, modifying the ResBlock structure in either the compression–reconstruction module or the subtraction module leads to significant performance changes.
TABLE V: Ablation Study on Hyperparameter Using the IRSTD-1K Dataset.
@@ IV-C7 The Effect of Pre-act ResBlock
IV-C8 Hyperparameter Discussion
The key hyperparameter in our LRRNet is the weight $\lambda$ in the loss function, which balances the optimization contributions from the segmentation and reconstruction branches. The impact of different $\lambda$ values on the final performance is shown in Tab. V. When $\lambda=0$, the model lacks constraints from the compression-reconstruction process, leading to degraded performance. On the other hand, a large $\lambda$ causes the reconstruction branch to dominate the optimization. An interesting observation during training is that when $\lambda$ is large, the segmentation loss remains nearly stagnant in the early stages and only begins to decrease when the reconstruction loss has been sufficiently minimized.
TABLE VI: Ablation Study on Model Size Using the IRSTD-1K Dataset.
@@ IV-C8 Hyperparameter Discussion
IV-C9 Model Size Discussion
Finally, we investigate the impact of architectural scale on model performance, focusing on the optimal configuration of channels, stages, and the number of ResBlocks per stage. As shown in Tab. VI, performance degrades when the number of channels is too small, as insufficient filters hinder the model’s ability to capture diverse image features. Conversely, blindly increasing the number of channels may also impair performance, as the additional capacity can lead to feature redundancy and diluted useful information. In terms of stage design, reducing the number of stages leads to out-of-memory (OOM) errors during training. This is because the self-attention module, applied at the lowest-resolution feature map, incurs a quadratic computational cost, which becomes prohibitive as the resolution increases. On the other hand, excessively increasing the number of stages results in over-parameterization, making the model prone to overfitting and degrading generalization. As illustrated in Fig. 3, due to the densely connected design of LRRNet, each decoder stage contains one more ResBlock than its corresponding encoder stage. For clarity, Tab. VI only reports the number of ResBlocks in each encoder stage under different configurations. It can be observed that the model achieves the best performance when each encoder stage contains two ResBlocks.
IV-D Limitations

### v3

@@ Abstract
This is the pre-acceptance version, to read the final version please go to IEEE Transactions on Geoscience and Remote Sensing on IEEE Xplore. Infrared small target detection (IRSTD) remains a long-standing challenge in complex backgrounds due to low signal-to-clutter ratios (SCR), diverse target morphologies, and the absence of distinctive visual cues. While recent deep learning approaches aim to learn discriminative representations, the intrinsic variability and weak priors of small targets often lead to unstable performance. In this paper, we propose a novel end-to-end IRSTD framework, termed LRRNet, which leverages the low-rank property of infrared image backgrounds. Inspired by the physical compressibility of cluttered scenes, our approach adopts a compression–reconstruction–subtraction (CRS) paradigm to directly model structure-aware low-rank background representations in the image domain, without relying on patch-based processing or explicit matrix decomposition. To the best of our knowledge, this is the first work to directly learn low-rank background structures using deep neural networks in an end-to-end manner. Extensive experiments on multiple public datasets demonstrate that LRRNet outperforms 38 state-of-the-art methods in terms of detection accuracy, robustness, and computational efficiency. Remarkably, it achieves real-time performance with an average speed of 82.34 FPS. Evaluations on the challenging NoisySIRST dataset further confirm the model’s resilience to sensor noise. Code has been available at LRRNet.
@@ I Introduction
The remainder of this paper is organized as follows. Section II provides a brief review of related work. Section III discusses several mathematical foundations of low-rank and sparse decomposition. Section IV presents the proposed method in detail. Experimental results and corresponding analyses are reported in Section V. Finally, Section VI concludes the paper.
@@ II-C Learned Image Compression Model
III Preliminaries
In this section, we discuss some mathematical theories of low-rank and sparse decomposition to further enhance the interpretability of our method.
Let us reformulate Eq. (1) into a decomposition form of the data matrix constructed from infrared patch-images, which is given as follows:
$\mathbf{D}=\mathbf{T}+\mathbf{B}+\mathbf{N},$ (2)
the explicit constraints of the low-rank and sparse decomposition are formulated as follows:
$\min\limits_{\mathbf{B},\mathbf{T}}rank(\mathbf{B})+\lambda\left\|\mathbf{T}\right\|_{0}\quad s.t.~\mathbf{D}=\mathbf{B}+\mathbf{T}\kern 5.0pt,\vskip-5.69046pt$ (3)
where $\lambda$ indicates a positive trade-off parameter, and $\left\|\cdot\right\|_{0}$} demotes the $l_{0}$-norm as the number of nonzero entries.
In this case, the deep unfolding framework involves alternately solving Eq. (3) via the ADMM algorithm, and subsequently transforming each iteration into a specific layer within the network. However, since the explicit low-rank constraint is imposed on patch-images rather than directly on the image domain, unfolding the model in the image domain may introduce inappropriate constraints. Once the patch-images are constructed, the resulting data matrix becomes extremely large. For example, a single $512\times 512$ image yields a $2500\times 2209$ matrix under the IPI [10] construction scheme. Moreover, as low-rankness is inherently a global property, it cannot be effectively enforced through conventional block-wise processing.
In contrast to deep unfolding methods, our core insight is that the low-rank property of image backgrounds is an inherent structural characteristic, and explicit matrix decomposition is merely one specific way to exploit this property—by no means the only one. Therefore, we interpret background low-rankness as a form of compressibility, aligning with the information-theoretic principle [20] that “learning is compression, and compression is learning.” Based on this perspective, we propose a compression–reconstruction–subtraction pipeline to enable small object detection.
IV Proposed Method
IV-A Overview
@@
IV-B Compression and Reconstruction Module
@@
Overview of the Proposed Compression and Reconstruction Module. The proposed compression-reconstruction module can be viewed as a specialized mapping $\mathcal{F}:\mathbf{F}_{I}\rightarrow\mathbf{F}_{B}$, where $\mathbf{F}_{I}\in\mathbb{R}^{W\times H\times C}$ denotes the feature map extracted from the original image, and $\mathbf{F}_{B}$ represents the corresponding reconstructed background feature map. Notably, $\mathbf{F}_{B}$ shares the same spatial and channel dimensions as $\mathbf{F}_{I}$. Given an input feature map $\mathbf{I}\in\mathbb{R}^{W\times H\times 1}$, we first apply a $3\times 3$ convolution to suppress local noise.
$\mathbf{F_{\mathit{I}}}=\mathrm{Conv}_{3\times 3}(\mathbf{I}).$ (4)
The processed features are then passed through successive stages of the densely connected U-Net. At the lowest resolution level, a self-attention mechanism [4] is employed to capture long-range dependencies between features.
@@ IV-D Loss Function
V Experiment
V-A Experimental Setup
V-A1 Datasets
@@ V-A2 Evaluation Metrics
where $N$ is the total number of samples, $T$, $P$, FPFP and $TP$ denote the number of ground truth, predicted positive pixels, the number of false positive, and the number of true positive respectively.
V-A3 Implementation Details
@@
V-B Comparison with State-of-the-Arts
@@
IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$ $\uparrow$ $F_{a}$ $\downarrow$ IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$ $\uparrow$ $F_{a}$ $\downarrow$ IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$ $\uparrow$ $F_{a}$ $\downarrow$
@@
ALCNet [18] TGRS’21 0.5404M 2.127G 62.05 59.58 92.19 31.56 $textemdash \textemdash \textemdash$ 61.13 60.61 96.51 29.09
FC3-Net [60] MM’22 6.896M 10.57G 64.98 63.59 92.93 15.73 $textemdash \textemdash \textemdash \textemdash \textemdash \textemdash \textemdash$
@@
TCI-Former [63] AAAI’24 3.66M 5.83G 70.14 67.69 96.31 14.81 $textemdash \textemdash \textemdash \textemdash \textemdash \textemdash \textemdash$
@@
$\text{Mrf}^{3}\text{Net}$ [19] TGRS’24 0.538M 33.2G 67.79 68.74 95.28 14.5 $textemdash \textemdash \textemdash$ 95.21 95.23 99.36 1.86
SCTransNet [65] TGRS’24 11.19M 67.4G 68.03 68.15 93.27 10.74 $textemdash \textemdash \textemdash$ 94.09 94.38 98.62 4.29
MLP-Net [66] TGRS’24 8.03M 8.714G 56.67 57.16 92.01 13.12 $textemdash \textemdash \textemdash$ 90.44 92.28 99.73 1.16
@@
CSRNet [9] TIP’24 0.4045M 121G 65.87 66.70 98.16 12.08 $textemdash \textemdash \textemdash \textemdash \textemdash \textemdash \textemdash$
OIPF-SCT [69] TAES’25 22.37M 80.97G 66.60 66.31 94.95 12.1 $textemdash \textemdash \textemdash$ 95.43 95.53 99.26 0.90
@@ V-B1 Quantitative Evaluation
TABLE III: Time Consumption of the Compared typical Methods (FPS)
IPI NIPPS RIPT PSTNN DNANet ALCNet RDIAN LRRNet
0.3 0.4 1.5 4.4 4.56 11.8 40 82.34
It is worth noting that our LRRNet achieves a throughput of 82.34 FPS on an NVIDIA 3080Ti GPU with an input resolution of $512\times 512$ and a batch size of 1. In contrast, as shown in Tab. III, methods that rely on learning object-specific features often suffer from significantly slower inference speeds [19]. For example, under the same conditions, DNANet [11] achieves only 4.56 FPS. Even RDIAN [30], which is specifically designed for lightweight architecture and high inference speed, still performs considerably slower than our LRRNet. This performance gap is primarily due to the use of complex module designs aimed at incorporating small-object information into deeper layers, which substantially increases computational overhead and slows down inference. Moreover, compared with traditional model-driven methods, LRRNet demonstrates significantly higher FPS. This is because LRRNet avoids time-consuming alternating iterations and instead performs target detection through a single forward pass, fully leveraging the highly parallel computing capabilities of GPUs.
@@ V-B2 Qualitative Evaluation
V-C Ablation Study
V-C1 The Effect of Densely Connection
@@
As shown in Table IV, the densely connected U-Net significantly outperforms the baseline: IoU increases by 1.90, nIoU by 1.8, $P_{d}$ by 1.68, and $F_{a}$ decreases by 11.29. This improvement stems from the limitation of standard skip connections, which, despite preserving information, suffer from feature confusion in deeper layers [43], hindering accurate background reconstruction. Moreover, deep equilibrium models [79] suggest that even identical modules can respond differently to various features. Dense long-range connections facilitate finer-grained feature propagation, enabling better preservation of subtle structures in cluttered scenes.
TABLE IV: Ablation Study on Compression–Reconstruction Module Using the IRSTD-1K Dataset.
@@
V-C2 The Effect of Long Skip Connections Without Attention
@@
In contrast, our design intentionally omits attention mechanisms along the skip paths. As shown in Table IV, this simple yet effective choice yields superior performance with statistically significant gains. We attribute this to the tendency of attention modules to overemphasize dominant structures, which, while beneficial for target enhancement, can hinder accurate low-rank background reconstruction. By avoiding such interference, our approach enables more precise modeling of the background and clearer separation between targets and clutter.
V-C3 The Effect of Self-Attention
@@
As shown in Table IV, introducing a single self-attention block significantly improves performance. However, adding multiple blocks leads to degradation. This can be attributed to the lack of inductive bias in self-attention [4]; when overused, it behaves like deep clustering—useful for compression but detrimental to reconstruction fidelity. Thus, a restrained application of self-attention is key to balancing compression efficiency with accurate reconstruction. Furthermore, we attempted to replace the self-attention mechanism with a tensor-based alternative, namely the Tensor Generation Module (TGM) and Tensor Reconstruction Module (TRM) proposed in [83], which perform explicit CP tensor decomposition. However, this substitution resulted in a significant drop in performance. We attribute this to the fact that explicit low-rank constraints in infrared imagery are more appropriately applied at the patch-image level rather than the image level. Direct decomposition in the full image domain imposes overly rigid constraints that hinder effective feature representation and reconstruction.
TABLE V: Ablation Study on Subtraction Module Using the IRSTD-1K Dataset.
@@
V-C4 The Effect of Subtraction Strategy
The subtraction strategy separates target-related information by subtracting the reconstructed background feature map from the corresponding feature map of the original image. If replaced with addition, the model effectively reverts to a conventional target feature learning approach. As shown in Tab. V, the subtraction strategy significantly enhances model performance.
V-C5 The Effect of Reconstruction Strategy
The reconstruction strategy is also a key component of LRRNet. On one hand, it guides the model to reconstruct the background; on the other, the use of MSE loss enhances the model’s robustness to Gaussian noise. As shown in Tab. V, the reconstruction strategy has a significant impact on overall performance.
V-C6 The Effect of ResBlock
The ResBlock plays an important role in effectively suppressing noise in the input image [55]. As shown in Tab. V, simply replacing it with a $3\times 3$ convolution results in a significant performance drop. However, increasing the number of ResBlocks excessively can also be detrimental. Although short skip connections help preserve features, combining them through nonlinear operations may introduce feature confusion. Given the small size of the targets, even minor disturbances can lead to noticeable performance degradation.
V-C7 The Effect of Pre-act ResBlock
The structure of the ResBlock also has a considerable impact on performance. In particular, the Pre-activation ResBlock demonstrates stronger robustness to noise compared to other variants [44], which is critical for both the compression–reconstruction and subtraction processes. As shown in Tab. IV and V, modifying the ResBlock structure in either the compression–reconstruction module or the subtraction module leads to significant performance changes.
TABLE VI: Ablation Study on Hyperparameter Using the IRSTD-1K Dataset.
@@ V-C7 The Effect of Pre-act ResBlock
V-C8 Hyperparameter Discussion
The key hyperparameter in our LRRNet is the weight $\lambda$ in the loss function, which balances the optimization contributions from the segmentation and reconstruction branches. The impact of different $\lambda$ values on the final performance is shown in Tab. VI. When $\lambda=0$, the model lacks constraints from the compression-reconstruction process, leading to degraded performance. On the other hand, a large $\lambda$ causes the reconstruction branch to dominate the optimization. An interesting observation during training is that when $\lambda$ is large, the segmentation loss remains nearly stagnant in the early stages and only begins to decrease when the reconstruction loss has been sufficiently minimized.
TABLE VII: Ablation Study on Model Size Using the IRSTD-1K Dataset.
@@ V-C8 Hyperparameter Discussion
V-C9 Model Size Discussion
Finally, we investigate the impact of architectural scale on model performance, focusing on the optimal configuration of channels, stages, and the number of ResBlocks per stage. As shown in Tab. VII, performance degrades when the number of channels is too small, as insufficient filters hinder the model’s ability to capture diverse image features. Conversely, blindly increasing the number of channels may also impair performance, as the additional capacity can lead to feature redundancy and diluted useful information. In terms of stage design, reducing the number of stages leads to out-of-memory (OOM) errors during training. This is because the self-attention module, applied at the lowest-resolution feature map, incurs a quadratic computational cost, which becomes prohibitive as the resolution increases. On the other hand, excessively increasing the number of stages results in over-parameterization, making the model prone to overfitting and degrading generalization. As illustrated in Fig. 3, due to the densely connected design of LRRNet, each decoder stage contains one more ResBlock than its corresponding encoder stage. For clarity, Tab. VII only reports the number of ResBlocks in each encoder stage under different configurations. It can be observed that the model achieves the best performance when each encoder stage contains two ResBlocks.
V-D Limitations